# Task-free Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v3

@@ Abstract
A fundamental objective in intelligent robotics is to move towards lifelong learning robot that can learn and adapt to unseen scenarios over time. However, continually learning new tasks would introduce catastrophic forgetting problems due to data distribution shifts. To mitigate this, we store a subset of data from previous tasks and utilize it in two manners: leveraging experience replay to retain learned skills and applying a novel Retrieval-based Local Adaptation technique to restore relevant knowledge. Since a lifelong learning robot must operate in task-free scenarios, where task IDs and even boundaries are not available, our method performs effectively without relying on such information. We also incorporate a selective weighting mechanism to focus on the most ”forgotten” skill segment, ensuring effective knowledge restoration. Experimental results across diverse manipulation tasks demonstrate that our framework provides a scalable paradigm for lifelong learning, enhancing robot performance in open-ended, task-free scenarios.
@@ 1 Introduction
Significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023). However, extending lifelong learning to robotics poses additional challenges, as robots must interact with the real world under sequential decision-making constraints. Besides, the high cost and complexity of physical interactions (Zhu et al., 2022; Du et al., 2023) limit data availability, emphasizing the need for more effective strategies to mitigate catastrophic forgetting and maintain performance over robots’ lifelong run (Thrun & Mitchell, 1995). Furthermore, in realistic and scalable scenarios, lifelong learning robots must operate in a task-free setting, where they are not provided with specific task IDs or boundaries for each new task. This further complicates the challenge, as robots must continually adapt to new tasks without prior knowledge of task distinctions.
In practical lifelong robot learning settings (Liu et al., 2024, 2023), a robot learns a series of tasks in a sequential manner. Lacking future data or environments, while also having only partial or no access to past data due to storage or security restrictions, leads to significant data distribution shifts and hinders stable lifelong learning. Although numerous approaches address this catastrophic forgetting problem (Wang et al., 2024), many rely on task boundaries or explicit task identifiers to consolidate knowledge, limiting their scalability in open-ended real-world scenarios (Koh et al., 2021).
@@
To tackle these issues, we propose a task-free memory-based approach for lifelong learning in robotic manipulation. Noteworthy, “task-free” does not mean the robot ignores the task context - it should be aware of the instructions combined with the observations to fulfill the job; rather, it means the proposed algorithm effectively handles multiple continually encountered tasks and integrates new knowledge and skills without relying on known task boundaries or explicit task IDs. The concept of “task-free” is properly defined in Section 3. Our method employs a compact storage memory $\mathcal{M}$ that holds a small set of previous tasks’ demonstrations. During training, we replay samples from $\mathcal{M}$ to preserve previously acquired knowledge and skills. Nonetheless, partial forgetting remains inevitable due to the multitasking nature of lifelong learning and limited access to past data. Inspired by human learning mechanisms — where we briefly revisit specific forgotten information to rapidly restore proficiency (Sara, 2000) — we then enable the robot to perform fast local adaptation just before policy deployment. Crucially, we use the same storage memory $\mathcal{M}$ for this adaptation, avoiding any additional storage burden.
Our system retrieves relevant demonstrations based on observation similarity (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) and selectively focuses on the most challenging segments where performance degrades—typically the final steps leading to task failure. This automatic selective weighting without relying on human intervention (Spencer et al., 2022; Mandlekar et al., 2020) enables the robots to efficiently restore lost knowledge and skills, facilitating stable, task-free lifelong robot learning. Our key contributions are summarized as:
@@
  • Task-free Retrieval-Based Local Adaptation: A novel local adaptation strategy that retrieves relevant past demonstrations from $\mathcal{M}$ to restore forgotten knowledge and skills, without requiring task IDs or boundaries.
@@
  • Selective Weighting Mechanism: An automatic weighting scheme emphasizing the most challenging segments within the retrieved demonstrations, optimizing real-time adaptation.
@@
  • A General Paradigm for Memory-Based Lifelong Robot Learning: Our approach serves as a plug-and-play solution, complementing existing memory-based lifelong learning algorithms and enabling skill restoration in sequences of open-ended robotics tasks.
@@ 2 Related Works
2.1 Lifelong Robot Learning
Robots operating in continually changing environments need the ability to learn and adapt on-the-fly (Thrun, 1995). In recent years, lifelong robot learning has been applied to SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022), navigation (Kim et al., 2024), and task and motion planning scenarios (Mendez-Mendez et al., 2023). In the context of lifelong learning from demonstrations, robots can 1) acquire skills from non-technical users (Grollman & Jenkins, 2007), or 2) adapt to user-specific task preferences — Chen et al. (2023) proposed a strategy mixture approach to efficiently model new incoming demonstrations, enhancing adaptability.
Furthermore, methods have been developed to improve manipulation capabilities over a robot’s lifespan. Some approaches leverage previous data to facilitate forward transfer but suffer from catastrophic forgetting (Xie & Finn, 2022). Others maintain an expandable skill set to accommodate an increasing number of manipulation tasks (Parakh et al., 2024), or continually update models of manipulable objects for effective reuse (Lu et al., 2022). Large language models have been utilized to improve knowledge transfer (Bärmann et al., 2023; Tziafas & Kasaei, 2024; Wang et al., 2023), and hypernetworks with neural ODEs have been employed to remember long trajectories (Auddy et al., 2023) incrementally. Additionally, (Yang et al., 2022) evaluates how typical supervised lifelong learning methods can be applied in reinforcement learning scenarios for robotic tasks.
@@ 2.1 Lifelong Robot Learning
To standardize the investigation of lifelong decision-making and bridge research gaps, Liu et al. (2024) introduced LIBERO, a benchmarking platform for lifelong robot manipulation where robots learn multiple atomic manipulation tasks sequentially. Recent works exploring lifelong robot learning based on it include (Liu et al., 2023), which assigns a specific task identity to each task; (Wan et al., 2024), which requires a pre-training phase to build an initial skill set before lifelong learning; and (Lee et al., 2024), tackles multi-stage tasks by incrementally learning skill prototypes for each subgoal, which introduces additional complexities in managing subgoal sequences. However, catastrophic forgetting for lifelong robot learning remains an open challenge, especially when task IDs and boundaries are not available.
2.2 Task-free Lifelong Learning
@@
Despite the success of lifelong learning under clearly labeled task sequences, a significant gap remains in algorithms that can operate independently of task boundaries or IDs during both training and inference, thus aligning more closely with realistic and scalable scenarios. Many approaches (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on specialized parameters via expanding network architectures. Meanwhile, researchers have tackled implicit task boundaries in regularization-based methods (Aljundi et al., 2019a; Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017) by consolidating knowledge upon detecting a loss “plateau”. Additionally, (Lässig et al., 2023) introduces a bio-inspired activity-regularization approach employing selective sparsity and recurrent lateral connections, effectively enabling task-free lifelong learning without explicit task boundaries.
Memory-based algorithms further mitigate forgetting by prioritizing informative samples (Sun et al., 2022), discarding less critical examples (Koh et al., 2021), refining decision boundaries (Shim et al., 2021), or enhancing gradient diversity (Aljundi et al., 2019b). Methods aiming to exploit replay buffer in online scenarios (Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success. However, these algorithms remain largely unexplored in robotic applications that entail sequential decision-making and real-world physical interactions.
2.3 Robot Learning with Adaptation
@@
Recent advances have shown robots adapting to dynamic environments, such as executing agile flight in strong winds (O’Connell et al., 2022), adapting quadruped locomotion through test-time search (Peng et al., 2020), and generalizing manipulation skills from limited data (Julian et al., 2020). To enable few-shot or one-shot adaptation, meta-learning has been extensively explored (Finn et al., 2017a) and successfully applied to robotics (Kaushik et al., 2020; Nagabandi et al., 2018; Finn et al., 2017b). However, meta-learning methods typically assume access to a full distribution of tasks during meta-training, with both training and testing performed on tasks sampled from this distribution. In contrast, our lifelong robot learning scenario operating sequentially lacks such access, presenting unique challenges of catastrophic forgetting.
@@
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and SentenceSimilarity (2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize a storage memory to store a subset of demonstrations gathered throughout the training process.
@@ 3 Preliminaries
Notably, multiple tasks may share overlapping distributions in either environmental settings or language descriptions. This natural setting closely mirrors real-world conditions, where it is difficult to determine which task generated a given scenario - tasks are not always divisible. This ambiguity underpins the proposed method’s task-free design, which emphasizes retrieving relevant information rather than relying on task boundaries or IDs.
@@
$\theta^{*}$ $=\arg\min_{\theta}\;\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\;g\sim G_{k}}$ (1)
$\quad\Biggl{[}\sum_{t=0}^{l_{k}}\mathcal{L}\bigl{(}\pi_{\theta}(o_{\leq t},g),\,a_{t}\bigr{)}\Biggr{]}$
@@
By optimizing this objective function, the policy effectively continues learning new knowledge and skills in its life span, without the need for task boundaries or IDs, thereby facilitating robust and adaptable task-free lifelong learning.
Figure 3: Data Retrieval. Episodic Memory $\mathcal{M}$ randomly stores a few demonstrations collected during lifelong learning. To retrieve a small number of demonstrations most similar to the current scenario, we compute a weighted distance (Eq (4.1)) using both image and language embeddings. In the storage memory $\mathcal{M}$, red and green circles denote relevant and irrelevant demonstrations, respectively, which include language descriptions, visual observations, expert actions, and joint and gripper states. The retrieved demonstrations are then used for Weighted Local Adaptation.
@@ 4 Retrieval-based Weighted Local Adaptation (RWLA) for Lifelong Robot Learning
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks. They produce consistent representations that are essential both for managing multiple tasks along the lifelong training and for retrieving relevant data to support our proposed local adaptation before policy deployment.
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in storage memory $\mathcal{M}$ (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024). Figure 2 illustrates the network architecture, and implementation details are provided in Appendix A.2.
@@ 4.1 Data Retrieval
The proposed task-free lifelong learning algorithm retrieves relevant demonstrations from $\mathcal{M}$ based on similarity to the deployment scenario $S_{deploy}$. Besides, due to the blurry task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, the retrieval process compares both visual inputs from the workspace camera (Du et al., 2023) and language descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings, following a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the language description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$, as illustrated from Figure 3.

### v4

# Task-agnostic Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

@@ Abstract
A fundamental objective in intelligent robotics is to move towards lifelong learning robots that can learn to manipulate in unseen scenarios over time. However, continually learning new tasks and manipulation skills from demonstration would introduce catastrophic forgetting due to data distribution shifts. To mitigate the problem, we store a subset of demonstrations from previous tasks and utilize them in two manners: leveraging experience replay to retain learned skills and applying a novel Retrieval-based Local Adaptation technique to recover relevant knowledge. Besides, task boundaries and IDs are unavailable in scalable, real-world settings, our method enables a lifelong learning robot to perform effectively without relying on such information. We also incorporate a selective weighting mechanism to focus on the most ”forgotten” action segment, ensuring effective skill recovery during adaptation. Experimental results across diverse manipulation tasks demonstrate that our framework provides a plug-and-play paradigm for lifelong learning, enhancing robot performance in open-ended, task-agnostic scenarios.
@@ 1 Introduction
Significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024; Cai & Müller, 2023; Gurbuz et al., 2024; Mai et al., 2021; Singh et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023; de Masson D’Autume et al., 2019; Biesialska et al., 2020). However, extending lifelong learning to robotics poses additional challenges, as robots must interact with the environment under sequential decision-making constraints. The high cost and complexity of physical interactions (Zhu et al., 2022; Du et al., 2023) limit the amount of available training data, making it critical to develop effective strategies to sustain robots’ long-term performance (Thrun & Mitchell, 1995). Additionally, in realistic and scalable scenarios, lifelong learning robots must operate in a task-agnostic setting, where they are not provided with specific task boundaries or IDs for each new task. This further complicates the challenge, as robots must continually learn without knowing task distinctions.
In practical lifelong robot learning settings (Liu et al., 2024; 2023) — distinct from offline pre/post-training on large cross-embodied datasets — robots learn tasks sequentially, with limited access to past data due to on-device storage, bandwidth, and privacy constraints. This process often faces significant task distribution shifts, leading to catastrophic forgetting. While many approaches have been proposed to address this challenge (Wang et al., 2024a), they often rely on task boundaries or IDs, limiting their scalability in open-ended real-world scenarios (Koh et al., 2021).
@@
To tackle these issues, we propose a task-agnostic, memory-based approach for lifelong learning of robotic manipulation skills from demonstrations. Noteworthy, “task-agnostic” does not mean the robot ignores the task context—it should be aware of language instructions and observations to fulfill the job; rather, it means the proposed algorithm effectively handles multiple continually encountered tasks and integrates new knowledge without relying on known task boundaries or IDs. Our method employs a compact storage memory $\mathcal{M}$ that holds a small set of previous tasks’ demonstrations. During training, we replay samples from $\mathcal{M}$ to preserve prior knowledge and skills. However, partial forgetting remains inevitable due to the multitasking nature of lifelong learning and the emphasis on training the current new task (Wang et al., 2024b). Therefore, instead of attempting to retain every detail throughout continual training, more effective mechanisms are needed to mitigate forgetting.
Human studies show that once knowledge is learned, even if it is forgotten over time, an efficient and targeted review can trigger memory retrieval and rapidly recover lost proficiency—often faster than learning for the first time (Sara, 2000; Roediger & Butler, 2011; Ebbinghaus, 2013; MacLeod, 1988). Inspired by the findings, we enable robots to perform fast local adaptation before policy deployment, allowing them to review and recover forgotten manipulation skills accumulated through lifelong learning. Crucially, we use the same storage memory $\mathcal{M}$ for adaptation, avoiding any additional storage burden. Our system retrieves relevant demonstrations based on contextual similarity (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) and automatically emphasizes the most failure-prone segments of each skill—typically critical steps where mistakes are likely to occur. This selective weighting, without requiring human intervention (Spencer et al., 2022; Mandlekar et al., 2020), promotes stable, task-agnostic lifelong learning. Our key contributions are summarized as:
@@
  • Task-agnostic Retrieval-Based Local Adaptation: A novel local adaptation strategy that retrieves relevant past demonstrations from $\mathcal{M}$ to recover forgotten skills, without requiring task boundaries or IDs.
@@
  • Selective Weighting Mechanism: An automated scheme that emphasizes the most challenging segments of retrieved demonstrations, optimizing real-time adaptation.
@@
  • A General Paradigm Solution: Our approach serves as a plug-and-play solution, complementing existing memory-based lifelong learning algorithms by enabling skill recovery in sequences of open-ended robotics tasks.
@@ 2 Related Works
Lifelong Robot Learning. Robots operating in continually changing environments need the ability to learn and adapt on-the-fly (Thrun, 1995; Grollman & Jenkins, 2007; Mendez-Mendez et al., 2023). In recent years, lifelong robot learning has been applied to SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022), navigation (Kim et al., 2024), and manipulation (Chen et al., 2023; Xie & Finn, 2022; Parakh et al., 2024; Lu et al., 2022; Auddy et al., 2023; Yang et al., 2022). Large language models have also been utilized to improve knowledge transfer (Bärmann et al., 2023; Tziafas & Kasaei, 2024; Wang et al., 2023).
@@
To standardize the investigation of lifelong decision-making and bridge research gaps, Liu et al. (2024) introduced LIBERO, a benchmarking platform for lifelong robot manipulation where robots learn multiple atomic manipulation tasks sequentially. Recent works exploring lifelong robot learning based on it include Liu et al. (2023), which assigns a specific task identity to each task; Wan et al. (2024), which requires a pre-training phase to build an initial skill set before lifelong learning; and Lee et al. (2024), tackles multi-stage tasks by incrementally learning skill prototypes for each subgoal, which introduces additional complexities in managing subgoal sequences. However, catastrophic forgetting for lifelong robot learning remains an open challenge, especially when task boundaries and IDs are not available.
@@
Task-agnostic Lifelong Learning. Despite the success of lifelong learning under clearly labeled task sequences, a significant gap remains in algorithms that can operate independently of task boundaries or IDs during both training and inference, thus aligning more closely with realistic and scalable scenarios. Many approaches (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on specialized parameters via expanding network architectures. Meanwhile, researchers have tackled implicit task boundaries in regularization-based methods by consolidating knowledge upon detecting a loss “plateau” (Aljundi et al., 2019a; Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017) or through a bio-inspired approach using selective sparsity and recurrent lateral connections (Lässig et al., 2023). Memory-based algorithms further mitigate forgetting by prioritizing informative samples (Sun et al., 2022), discarding less critical examples (Koh et al., 2021), refining decision boundaries (Shim et al., 2021), or enhancing gradient diversity (Aljundi et al., 2019b). Methods aiming to exploit the replay buffer in online scenarios (Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success. However, these algorithms remain largely unexplored in robotic applications that entail sequential decision-making and real-world physical interactions.
@@
Robot Learning with Adaptation. Recent advances have shown robots adapting to dynamic environments, such as executing agile flight in strong winds (O’Connell et al., 2022), adapting quadruped locomotion through test-time search (Peng et al., 2020), and generalizing manipulation skills from limited data (Julian et al., 2020; Memmel et al., 2024; Lin et al., 2024). To enable few-shot or one-shot adaptation, meta-learning has been extensively explored (Finn et al., 2017a) and successfully applied to robotics (Kaushik et al., 2020; Nagabandi et al., 2018; Finn et al., 2017b; Schmied et al., 2023). However, meta-learning methods typically assume access to a full distribution of tasks during meta-training, with both training and testing performed on tasks sampled from this distribution. In contrast, our lifelong robot learning scenario operating sequentially lacks such access, presenting unique challenges of catastrophic forgetting.
@@
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and (SentenceSimilarity, 2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize a storage memory to store a subset of demonstrations gathered throughout the training process.
@@ 3 Preliminaries
Notably, multiple tasks may share overlapping distributions in either environmental settings or language descriptions. This natural setting closely mirrors real-world conditions, where it is difficult to determine which task generated a given scenario - tasks are not always divisible. This ambiguity underpins the proposed method’s task-agnostic design, which emphasizes retrieving relevant information rather than relying on task boundaries or IDs.
@@
$\theta^{*}=\arg\min_{\theta}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\,g\sim G_{k}}\left[\sum_{t=0}^{l_{k}}\mathcal{L}\left(\pi_{\theta}(o_{\leq t},g),\,a_{t}\right)\right],$ (1)
@@
By optimizing this objective function, the policy effectively continues learning new knowledge and skills in its life span, without the need for task boundaries or IDs, thereby facilitating robust and adaptable task-agnostic lifelong learning.
Algorithm 1 RWLA for Task-agnostic Lifelong Robot Learning
Lifelong Learning Phase:
  1. 1.
  2. 2.
    $K\in\{1,2,\dots,T\}$
    1. (a)
      Train $\theta$ on $\mathcal{D}_{K}\cup\mathcal{M}$ using Eq 1
    2. (b)
Reviewing Phase:
  1. 1.
  2. 2.
    Retrieve $\tilde{N}$ demonstrations from $\mathcal{M}$ based on embedding distance using Eq 2 (Section 4.1)
  3. 3.
    Compute $w_{t,n}$ based on selective weighting (Section 4.2.1)
  4. 4.
    $\theta^{\prime}\leftarrow$ Locally adapt $\theta$ using Eq 3 as skill recovery within limited epochs (Section 4.2.2)
Testing Phase: Test $\theta^{\prime}$ in $\mathcal{S}_{deploy}$
@@ 4 Retrieval-based Weighted Local Adaptation (RWLA) for Lifelong Robot Learning
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks. They produce consistent representations that are essential both for managing multiple tasks throughout lifelong training and for retrieving relevant data to support our proposed local adaptation before policy deployment.
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in storage memory $\mathcal{M}$ (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024a). Figure 2 illustrates the network architecture, and implementation details are provided in Appendix A.2.
@@ 4.1 Data Retrieval
The proposed task-agnostic lifelong learning algorithm retrieves relevant demonstrations from $\mathcal{M}$ based on similarity to the deployment scenario $S_{deploy}$. Besides, due to the blurry task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, the retrieval process compares both visual inputs from the workspace camera (Du et al., 2023) and language descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings, following a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$ (2)
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the language description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$, as illustrated in Figure 3.